OpenStudy (anonymous):

hi can u help in q4 X,Y,Z values and if possible, with Q1, and q2 part to find "gm" thank you @halfwebprince @radar @malihe @mmpt @monishvster @wmuldoon26 @sagar93000 @dan815 @DejohnetteW @hiden @jutt @ka.shu @batenhafiz @needhelpquick111 @Nurali

OpenStudy (anonymous):

dude Q4 is still unsolved

OpenStudy (anonymous):

q4 half solved. only xyz values are remaining?? I thing we need to apply concepts of control system...

OpenStudy (anonymous):

what's the formula for q1 part 5??

OpenStudy (anonymous):

@monishvster wat r ur paramters for that question

OpenStudy (anonymous):

V0=40V, R1=100Ω, R2=100Ω, R3=200Ω, L=1H, and C=100μF

OpenStudy (anonymous):

if anyone has more tries left check them x=LC y=R2C z=1

OpenStudy (anonymous):

x=LC is wrong z=1 is wrong

OpenStudy (anonymous):

is y correct??

OpenStudy (anonymous):

y I don't know

OpenStudy (anonymous):

then don't waste ur tries. i'll try again

OpenStudy (anonymous):

go to page 4-5 and solve accordingly . r3 will be eliminated due to current flowing through short circuit path.

OpenStudy (anonymous):

X=R1LC Y=L+R1R2C Z=R1+R2

OpenStudy (anonymous):

@kartkn can u plz post the formula of Q4 first 4 problems

OpenStudy (anonymous):

no formulas

OpenStudy (anonymous):

so far 85% course done guys.. .

OpenStudy (anonymous):

Q4: 1.vc(t<0)=v0*(R2+R3)/(R1+R2+R3) 2.ic(t<0)=v0/(R1+R2+R3) 3.vc(t>>0)=v0*R2/(R1+R2) 4.ic(t>>0)=v0/(R1+R2)

OpenStudy (anonymous):

is q4 correct? thanks for your help ! ;)